OpenStudy (anonymous):

what is a10 for the sequence an = 1 + 3 (n - 1)?

OpenStudy (mathmath333):

just put the value of n=10 in the formula \(\large\tt \color{black}{an = 1 + 3 (n - 1)}\) \(\large\tt \color{black}{a10 = 1 + 3 (10 - 1)}\) \(\large\tt \color{black}{a10 = 1 + 3 (9)}\) \(\large\tt \color{black}{a10 = 1 + 27}\) \(\large\tt \color{black}{a10 = 28}\)

OpenStudy (anonymous):

Thank you.

OpenStudy (mathmath333):

welcome